# Time-Correlated Video Bridge Matching

### v1

@@ Abstract
Diffusion models excel in noise-to-data generation tasks, providing a mapping from a Gaussian distribution to a more complex data distribution. However they struggle to model translations between complex distributions, limiting their effectiveness in data-to-data tasks. While Bridge Matching (BM) models address this by finding the translation between data distributions, their application to time-correlated data sequences remains unexplored. This is a critical limitation for video generation and manipulation tasks, where maintaining temporal coherence is particularly important. To address this gap, we propose Time-Correlated Video Bridge Matching (TCVBM), a framework that extends BM to time-correlated data sequences in the video domain. TCVBM explicitly models inter-sequence dependencies within the diffusion bridge, directly incorporating temporal correlations into the sampling process. We compare our approach to classical methods based on bridge matching and diffusion models for three video-related tasks: frame interpolation, image-to-video generation, and video super-resolution. TCVBM achieves superior performance across multiple quantitative metrics, demonstrating enhanced generation quality and reconstruction fidelity.
@@ 1 INTRODUCTION
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have emerged as a powerful paradigm for generative modeling, achieving remarkable results in high-fidelity data synthesis (Saharia et al., 2022; Rombach et al., 2022; Arkhipkin et al., 2024; Labs, 2024). By iteratively denoising samples from a Gaussian distribution, these models excel at producing diverse and realistic outputs. However, despite their widespread adoption, diffusion models exhibit a critical limitation: they struggle to model translations between complex-structured data distributions. This shortcoming hinders their effectiveness in data-to-data tasks, where smooth bridging of the distributions is essential.
In contrast, Bridge Matching (BM) offers a principled solution to this problem by explicitly constructing a translation between arbitrary data distributions (Peluchetti, 2023b, a; Liu et al., 2022; Zhou et al., 2023). These methods learn vector field that connects source and target distributions, demonstrating strong performance in image-to-image tasks  (Shi et al., 2023; Liu et al., 2023). However, the translation between independent data sequences, the components of which are time-correlated, remains unaddressed. At the same time, this is exactly what video data is, where one data sample is a sequence of correlated frames. While existing methods assume to operate with video data samples without considering their internal structure  (Wang et al., 2025b), this omission can lead to a decrease in temporal consistency in the generated video.
Contribution. To address these issues, we propose Time-Correlated Video Bridge Matching (TCVBM), a novel framework that extends Bridge Matching to time-dependent video data. Unlike prior work, TCVBM explicitly takes into account inter-sequence dependencies, ensuring faithful translation between source and target distributions (Figure 1). We evaluate TCVBM on three video-related tasks: frame interpolation, image-to-video generation, and video super resolution. A comparison with classical approaches that do not take into account the correlation between frames, such as DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching, demonstrates that TCVBM provides better temporal consistency and reconstruction quality.
@@ 2.1 Bridge Models
Despite the widespread adoption of diffusion models in generative tasks (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b), the Gaussian noise commonly used to initialize the generation process lacks meaningful structural information about the underlying data distribution. However, by matching the velocity fields with pre-defined transport maps, it is possible to construct powerful generative models that rival diffusion-based approaches (Lipman et al., 2023). Furthermore, Diffusion Bridge Models have been shown to provide a more flexible framework for generative tasks compared to standard diffusion processes (Bortoli et al., 2021; Liu et al., 2022). Particularly, bridge models have shown their superiority over diffusion baselines for image restoration (Delbracio and Milanfar, 2024), image translation (Zhou et al., 2024a), and image reconstruction (Liu et al., 2023). Despite this, the use of Bridge Matching to work with complex distributions of sequential data, the components of which are correlated, remains a practically unexplored topic. A recent attempt to extend Bridge Matching to image-to-video generation, overlooked the intrinsic temporal dependencies between video frames (Wang et al., 2025b). In our solution, we address this problem and design our own interpolant that explicitly accounts for the assumption of linear correlations among the components of the random vectors that represent video data samples.
@@ 2.2 Video Generation and Manipulation
Here, we provide a brief overview of the methods for some video-related tasks. Without loss of generality, we evaluate the applicability of our approach to the video domain using the following three tasks:
@@
aims to synthesize middle frames between two input images while ensuring temporal smoothness and content consistency. Traditional approaches often rely on optical flow estimation between keyframes (Niklaus and Liu, 2020; Lee et al., 2020; Park et al., 2021; Huang et al., 2022) or leverage convolutional features and attention maps for middle frame generation (Kalluri et al., 2023; Shi et al., 2022; Reda et al., 2022). Diffusion models were first used for video interpolation using bidirectional masking to synthesize middle frames (Voleti et al., 2022). Subsequent advances expanded this direction, including conditional diffusion-based generation (Danier et al., 2024), multi-stage cascaded refinement (Jain et al., 2024), adapted image-to-video diffusion models (Wang et al., 2025a), and techniques for large-motion generation (Shen et al., 2024). However, these methods do not explicitly model inter-frame correlations in video sequences. Similarly, event-based approaches (Chen et al., 2025; Zhang et al., 2025) incorporate additional motion cues but still rely on standard diffusion processes without accounting for temporal correlations.
@@
is a specific type of video generation, where an image is used as the input for a neural network, which then creates a video with consistent and semantically accurate movement. Diffusion models and recent advancements in generative architectures have taken this task to a new level, providing high-quality results (Zhang et al., 2023; Xing et al., 2023; Shi et al., 2024; Guo et al., 2023, 2024; Arkhipkin et al., 2025). However, the fundamental limitation of the diffusion process on the noise-to-data setup posed a risk of losing essential information that must be transferred from the image to the animated video. Different approaches tried to solve this without going beyond the diffusion process (Ren et al., 2024; Wu et al., 2023), until FrameBridge (Wang et al., 2025b) proposed an approach for image-to-video generation based on Bridge Matching. FrameBridge uses the input image as a prior instead of the Gaussian distribution used in the diffusion algorithm. Although this makes it reasonable to consider the image-to-video generation as a data-to-data task, this approach does not take into account the specifics of the internal video data structure, in fact adapting bridge matching algorithm from image to video domain. In this paper, we propose a method that explicitly assumes the correlation of frames with each other and uses this to obtain samples from bridge distribution.
@@
focuses on reconstructing high-resolution video frames from their degraded low-resolution versions. The application of diffusion models to this task is motivated by their strong generative prior, which is capable of synthesizing realistic details to overcome complex degradations. A central challenge is reconciling the stochastic nature of the diffusion process with the stringent requirement for temporal coherence across video frames. To resolve this, methods such as Upscale-A-Video (Zhou et al., 2024b) and MGLD (Yang et al., 2024) introduce explicit spatiotemporal constraints, using techniques like temporal layer integration and motion-guided loss optimization. At the same time, bridge-matching methods have been applied for super resolution in the image domain (Liu et al., 2023; Gushchin et al., 2025). In this work, we expand the application of these methods to video super resolution and propose a new approach that explicitly accounts for temporal coherence between video frames.
@@ 3 BACKGROUND ON BRIDGE MATCHING
We briefly review the Bridge Matching framework (Peluchetti, 2023b, a; Liu et al., 2022; Shi et al., 2023), which constructs diffusion processes for data translation, given a distribution of clean data $p(\mathbf{x}_{0})$ and corrupted data $p(\mathbf{x}_{T})$ on $\mathbb{R}^{D}$. The goal is to model a stochastic process that transitions from $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ to $\mathbf{x}_{T}\sim p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, while incorporating a prior dynamics.
@@ 4.1 Time-Correlated Prior Process
$\boldsymbol{\Sigma}_{t}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t})\in\mathbb{R}^{N\times D}$
denotes applying $\boldsymbol{\Sigma}_{t}^{-1}\in\mathbb{R}^{N\times N}$ independently to each column of $\mathbf{X}_{t}-\boldsymbol{\mu}_{t}$. Further, unless otherwise stated, we will assume a time-independent noise scale $g(t)=\sqrt{\epsilon}$ for simplicity.
@@
$q(\mathbf{X}_{t}|\mathbf{X}_{0})=\mathcal{N}(\mathbf{X}_{t}\mid\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}),\boldsymbol{\Sigma}_{t|0}),$ (5)
@@
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})$ $=e^{\mathbf{A}t}\mathbf{X}_{0}+\left(e^{\mathbf{A}t}-I\right)\mathbf{A}^{-1}\mathbf{b},$ (6)
$\boldsymbol{\Sigma}_{t|0}$ $=\epsilon\,\frac{e^{2\mathbf{A}t}-I}{2}\mathbf{A}^{-1}.$ (7)
@@
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}|\mathbf{X}_{0})=-\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})).$ (8)
@@
$q(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{t^{\prime}})=\mathcal{N}(\mathbf{X}_{t}\mid\boldsymbol{\mu}_{t|0,t^{\prime}},\boldsymbol{\Sigma}_{t|0,t^{\prime}}),$ (9)
@@
$\boldsymbol{\mu}_{t|0,t^{\prime}}$ $=\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})+\boldsymbol{\Sigma}_{t|0}\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}(\mathbf{X}_{t^{\prime}}-\boldsymbol{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})),$ (10)
$\boldsymbol{\Sigma}_{t|0,t^{\prime}}$ $=\boldsymbol{\Sigma}_{t|0}-\boldsymbol{\Sigma}_{t|0}\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}\boldsymbol{\Sigma}_{t|0}.$ (11)
@@ 4.2 Time-Correlated Video Bridge Matching
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$ (12)
@@
The minimizer $v^{*}(\mathbf{X}_{t},t)$ of the objective (12) can be expressed as:
$v^{*}(\mathbf{X}_{t},t)=-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right),$
@@
We parameterize the predictor $\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)$ with a neural network and train it using the regression loss in (13). The training procedure is summarized in Algorithm 1.
@@
0: data from coupling $p_{0}(\mathbf{X}_{0})p_{T}(\mathbf{X}_{T}|\mathbf{X}_{0})$ and coefficients $\mathbf{A}$, $\mathbf{b}$ and $\epsilon$ for prior (1).
@@
3:  $\mathbf{X}_{t}\sim q\bigl(\mathbf{X}_{t}\mid\mathbf{X}_{0},\mathbf{X}_{T}\bigr)$ (9)
4:  Take gradient descent step on $\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)$ (13)
@@
4:  Sample $\mathbf{X}_{t_{n-1}}\sim p(\mathbf{X}_{t_{n-1}}\mid\widehat{\mathbf{X}}_{0},\mathbf{X}_{t_{n}})$ using (9)
@@ 4.3 Choice of the Prior Process for Video Manipulation Tasks
Figure 2: Frame interpolation results. The first and last frames are fixed and come from the dataset. The task is to generate the intermediate 88 frames. TCVBM provides a smoother and more consistent transition from frame to frame and a better detalization.

### v2

@@ Abstract
Diffusion models excel in noise-to-data generation tasks, providing a mapping from a Gaussian distribution to a more complex data distribution. However they struggle to model translations between complex distributions, limiting their effectiveness in data-to-data tasks. While Bridge Matching models address this by finding the translation between data distributions, their application to time-correlated data sequences remains unexplored. This is a critical limitation for video generation and manipulation tasks, where maintaining temporal coherence is particularly important. To address this gap, we propose Time-Correlated Video Bridge Matching (TCVBM), a framework that extends BM to time-correlated data sequences in the video domain. TCVBM explicitly models inter-sequence dependencies within the diffusion bridge, directly incorporating temporal correlations into the sampling process. We compare our approach to classical methods based on bridge matching and diffusion models for three video-related tasks: frame interpolation, image-to-video generation, and video super-resolution. TCVBM achieves superior performance across multiple quantitative metrics, demonstrating enhanced generation quality and reconstruction fidelity.
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have emerged as a powerful paradigm for generative modeling, achieving remarkable results in high-fidelity data synthesis (Saharia et al., 2022; Rombach et al., 2022; Arkhipkin et al., 2024; Labs, 2024; Arkhipkin et al., 2025a). By iteratively denoising samples from a Gaussian distribution, these models excel at producing diverse and realistic outputs. However, despite their widespread adoption, diffusion models exhibit a critical limitation: they struggle to model translations between complex-structured data distributions. This shortcoming hinders their effectiveness in data-to-data tasks, where smooth bridging of the distributions is essential.
In contrast, Bridge Matching (BM) offers a principled solution to this problem by explicitly constructing a translation between arbitrary data distributions (Peluchetti, 2023b; a; Liu et al., 2022; Zhou et al., 2023). These methods learn vector field that connects source and target distributions, demonstrating strong performance in image-to-image tasks  (Shi et al., 2023; Liu et al., 2023). However, the translation between independent data sequences, the components of which are time-correlated, remains unaddressed. At the same time, this is exactly what video data is, where one data sample is a sequence of correlated frames. While existing methods assume to operate with video data samples without considering their internal structure  (Wang et al., 2025b), this omission can lead to a decrease in temporal consistency in the generated video.
Contribution. To address these issues, we propose Time Correlated Video Bridge Matching (TCVBM), a novel framework that extends Bridge Matching to time-dependent video data. Unlike prior work, TCVBM explicitly takes into account inter-sequence dependencies, ensuring faithful translation between source and target distributions (Figure 1). We evaluate TCVBM on three video-related tasks: frame interpolation, image-to-video generation, and video super resolution. A comparison with classical approaches that do not take into account the correlation between frames, such as DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching, demonstrates that TCVBM provides better temporal consistency and reconstruction quality.
@@ 2.1 Bridge Models
Despite the success of diffusion models in generative tasks (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b), their reliance on Gaussian noise as a prior lacks meaningful structural information about the data. In contrast, models that match velocity fields using pre-defined transport maps can achieve competitive performance (Lipman et al., 2023). Bridge Matching offers a particularly flexible framework, outperforming standard diffusion for tasks like image restoration, translation, and reconstruction (Delbracio and Milanfar, 2024; Zhou et al., 2024a; Liu et al., 2023). However, applying Bridge Matching to correlated sequential data, such as video, remains largely unexplored. A recent extension to image-to-video generation (Wang et al., 2025b) overlooked inherent temporal dependencies. Our approach addresses this by designing an interpolant that explicitly models the linear correlations between video frames.
2.2 Temporal Modeling for Video Data
Advances in video generation often adapt pre-trained image models by adding temporal modules like 3D convolutions or attention layers (Ho et al., 2022; Blattmann et al., 2023; Arkhipkin et al., 2023). This architectural specialization continues with Diffusion Transformer-based video models (Chen et al., 2023; Ma et al., 2025). A key challenge is the computational cost of attention, addressed by techniques such as sparse attention and adaptive masking (Zhang et al., 2025b; Xi et al., 2025; Mikhailov et al., 2025). However, modeling temporal dependencies within the generative dynamics and SDE prior structure remains unexplored. These prior works are orthogonal to our contribution and compatible with our approach.
@@ 2.3 Video Generation and Manipulation
aims to synthesize middle frames from two inputs, ensuring smoothness and consistency. Traditional methods use optical flow (Niklaus and Liu, 2020; Lee et al., 2020; Park et al., 2021; Huang et al., 2022) or convolutional features with attention (Kalluri et al., 2023; Shi et al., 2022; Reda et al., 2022). Diffusion-based interpolation began with bidirectional masking (Voleti et al., 2022) and advanced through conditional generation (Danier et al., 2024), cascaded refinement (Jain et al., 2024), adapted image-to-video models (Wang et al., 2025a), and large-motion techniques (Shen et al., 2024). However, these methods lack explicit modeling of inter-frame correlations. Event-based approaches (Chen et al., 2025; Zhang et al., 2025a) add motion cues but also use standard diffusion without capturing temporal dependencies.
@@
creates a video from an input image, requiring consistent and accurate motion. Diffusion models have significantly advanced this field, producing high-quality results (Zhang et al., 2023; Xing et al., 2023; Shi et al., 2024; Guo et al., 2023; 2024; Arkhipkin et al., 2025b; a). However, the standard noise-to-data diffusion process risks losing essential information from the input image. While some approaches address this within the diffusion framework (Ren et al., 2024; Wu et al., 2023), FrameBridge (Wang et al., 2025b) reformulates the task using Bridge Matching, treating the input image as a prior instead of Gaussian noise. This frames generation as a data-to-data task but does not account for the inherent structure and correlations between video frames. Our proposed method explicitly models these inter-frame correlations to sample from the bridge distribution.
@@
reconstructs high-resolution videos from low-resolution inputs. Diffusion models are applied for their strong generative prior, which synthesizes realistic details to overcome degradation. A key challenge is ensuring temporal coherence within the inherently stochastic diffusion process. Recent methods address this by introducing explicit spatiotemporal constraints, such as temporal layer integration and motion-guided losses (Zhou et al., 2024b; Yang et al., 2024). Meanwhile, bridge-matching methods have shown promise for image super-resolution (Liu et al., 2023; Gushchin et al., 2025). In this work, we extend bridge matching to video super resolution, proposing a novel approach that explicitly models temporal coherence between frames.
@@ 3 Background on Bridge Matching
We briefly review the Bridge Matching framework (Peluchetti, 2023b; a; Liu et al., 2022; Shi et al., 2023), which constructs diffusion processes for data translation, given a distribution of clean data $p(\mathbf{x}_{0})$ and corrupted data $p(\mathbf{x}_{T})$ on $\mathbb{R}^{D}$. The goal is to model a stochastic process that transitions from $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ to $\mathbf{x}_{T}\sim p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, while incorporating a prior dynamics.
@@ 4.1 Time-Correlated Prior Process
$\bm{\Sigma}_{t}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t})\in\mathbb{R}^{N\times D}$
denotes applying $\bm{\Sigma}_{t}^{-1}\in\mathbb{R}^{N\times N}$ independently to each column of $\mathbf{X}_{t}-\bm{\mu}_{t}$. Further, unless otherwise stated, we will assume a time-independent noise scale $g(t)=\sqrt{\epsilon}$ for simplicity.
@@
$q(\mathbf{X}_{t}|\mathbf{X}_{0})=\mathcal{N}(\mathbf{X}_{t}\mid\bm{\mu}_{t|0}(\mathbf{X}_{0}),\bm{\Sigma}_{t|0}),$ (5)
@@
$\bm{\mu}_{t|0}(\mathbf{X}_{0})$ $=e^{\mathbf{A}t}\mathbf{X}_{0}+\left(e^{\mathbf{A}t}-I\right)\mathbf{A}^{-1}\mathbf{b},$ (6)
$\bm{\Sigma}_{t|0}$ $=\epsilon\,\frac{e^{2\mathbf{A}t}-I}{2}\mathbf{A}^{-1}.$ (7)
@@
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}|\mathbf{X}_{0})=-\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0})).$ (8)
@@
$q(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{t^{\prime}})=\mathcal{N}(\mathbf{X}_{t}\mid\bm{\mu}_{t|0,t^{\prime}},\bm{\Sigma}_{t|0,t^{\prime}}),$ (9)
@@
$\bm{\mu}_{t|0,t^{\prime}}$ $=\bm{\mu}_{t|0}(\mathbf{X}_{0})+\bm{\Sigma}_{t|0}\bm{\Sigma}_{t^{\prime}|0}^{-1}(\mathbf{X}_{t^{\prime}}-\bm{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})),$ (10)
$\bm{\Sigma}_{t|0,t^{\prime}}$ $=\bm{\Sigma}_{t|0}-\bm{\Sigma}_{t|0}\bm{\Sigma}_{t^{\prime}|0}^{-1}\bm{\Sigma}_{t|0}.$ (11)
@@ 4.2 Time-Correlated Video Bridge Matching
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$ (12)
@@
The minimizer $v^{*}(\mathbf{X}_{t},t)$ of the objective equation 12 can be expressed as:
$v^{*}(\mathbf{X}_{t},t)=-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right),$
@@
We parameterize the predictor $\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)$ with a neural network and train it using the regression loss in equation 13. The training procedure is summarized in Algorithm 1.
@@
0: data from coupling $p_{0}(\mathbf{X}_{0})p_{T}(\mathbf{X}_{T}|\mathbf{X}_{0})$ and coefficients $\mathbf{A}$, $\mathbf{b}$ and $\epsilon$ for prior equation 1.
@@
3:  $\mathbf{X}_{t}\sim q\bigl(\mathbf{X}_{t}\mid\mathbf{X}_{0},\mathbf{X}_{T}\bigr)$ equation 9
4:  Take gradient descent step on $\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)$ equation 13
@@
4:  Sample $\mathbf{X}_{t_{n-1}}\sim p(\mathbf{X}_{t_{n-1}}\mid\widehat{\mathbf{X}}_{0},\mathbf{X}_{t_{n}})$ using equation 9
@@ 4.3 Choice of the Prior Process for Video Manipulation Tasks
5 Experiments
5.1 Experiments Roadmap
We compare TCVBM with DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching (BM) (Ibe, 2013, Chapter 9). Experiments use the MovingMNIST dataset (Srivastava et al., 2015), which contains $10,000$ video sequences each $20$ frames long and showing $2$ moving digits in a $64\times 64$ resolution. Implementation details can be found in the Appendix B. Using each approach, we train three models for the following tasks:
Frame Interpolation.
The network inputs a sequence of $10$ frames, with the first and last frames fixed from the dataset. It generates $8$ middle frames, which are compared to the ground truth. DDPM and DDIM generate these frames from random noise. For BM and TCVBM, we tested initializing the middle frames via linear interpolation or Gaussian noise from $\mathcal{N}(\textbf{0},\mathbf{I})$; noise yielded better results and is used further. Details can be found in the Appendix D.1.
Image-to-Video Generation.
The task is to generate the remaining $9$ frames based on the first frame. We provide the first frame as input data, along with $9$ Gaussian samples from $\mathcal{N}(\textbf{0},\mathbf{I})$, or $9$ copies of the first frame. The second option demonstrated the best results for all methods, so further we consider this initialization variant. For more information, please see the Appendix D.2.
(a) Frame interpolation. The first and last frames are fixed and come from the dataset. The task is to generate the intermediate 88 frames. TCVBM provides a smoother and more consistent transition from frame to frame and a better detalization.
(b) Image-to-Video generation. Based on the first frame, the method generate the next 99 frames. As can be seen, TCVBM makes it possible to achieve better appearance of the generations and a lower quality degradation with increasing sequence length.
(c) Video super resolution. The resolution is increased from $16\times 16$ to $64\times 64$. As can be seen, TCVBM outperforms other methods.
Figure 2: The results on the MovingMNIST dataset Srivastava et al. (2015).
Video Super Resolution.
The network inputs $10$ low-resolution frames and outputs $10$ frames at the dataset resolution. We evaluated starting resolutions of $16\times 16$ and $32\times 32$; as the simpler task showed no significant advantage for our method at $32\times 32$, we report results for $16\times 16$. We also tested initialization with low-resolution data alone versus data concatenated with Gaussian noise; both performed similarly, and we proceed with the noise-free version. See details in Appendix D.3.
We present results for BM and TCVBM with $\epsilon=0.1$ and $\alpha=1$ for TCVBM (i.e., $\widetilde{\mathbf{A}}=\mathbf{A}$, $\widetilde{\mathbf{b}}=\mathbf{b}$) for all three tasks. Experiments optimizing hyperparameters $\epsilon$ and $\alpha$ are in Appendix E. Computational complexity analysis is in Appendix F. We also explored a dynamic correlation approach where $\alpha_{t}$ increases as $t\to 0$, but found no advantage over constant $\alpha$ (see Appendix G).
5.2 Results
Qualitative Comparison.
As shown in Figure 2, TCVBM yields more consistent and accurate results than other methods. By modeling mutual correlation between frames, it ensures smoother transitions and better information transfer, improving overall stability. More examples can be found in the Appendix H.